if the earth be one half of its present distance from the sun the number of the days in the year will be nearly

Grade:12th pass

1 Answers

Arun
6 years ago
The ratio of the time periods would be {new is say T1, actual is say T2}
Note that T is directly proportional to √(R3)
T1/T2 = √([R/2]3)/√(R3) = 1/(2√2)
Hence T1 = T2/2√2 = 365/2√2
Which is approx 129 days.

Regards
Arun (askIITians forum expert)